### 04 Dec 2023 v1

Abstract
Past studies have empirically demonstrated a surprising agreement between gravitational waveforms computed using adiabatic-driven-inspiral point-particle black hole perturbation theory (ppBHPT) and numerical relativity (NR) following a straightforward calibration step, sometimes referred to as - scaling. Specifically focusing on the quadrupole mode, this calibration technique necessitates only two time-independent parameters to scale the overall amplitude and time coordinate. In this article, part of a special issue, we investigate this scaling for non-spinning binaries at the equal mass limit. Even without calibration, NR and ppBHPT waveforms exhibit an unexpected degree of similarity after accounting for different mass scale definitions. Post-calibration, good agreement between ppBHPT and NR waveforms extends nearly up to the point of the merger. We also assess the breakdown of the time-independent assumption of the scaling parameters, shedding light on current limitations and suggesting potential generalizations for the - scaling technique.
keywords:
[…]
1 Introduction
Simulating binary black hole (BBH) mergers and understanding the morphology of the resultant gravitational waveforms enables us to develop tools for the rapid characterization of detected gravitational wave (GW) signals Blackman et al. (2015, 2017a, 2017b); Varma et al. (2019a, b); Islam et al. (2021); Bohé et al. (2017); Cotesta et al. (2018, 2020); Pan et al. (2014); Babak et al. (2017); Husa et al. (2016); Khan et al. (2016); London et al. (2018); Khan et al. (2019). An accurate description of BBH mergers requires numerically solving the Einstein equation for the two-body problem. This approach is known as numerical relativity (NR) Mroue et al. (2013); Boyle et al. (2019); Healy et al. (2017, 2019); Healy and Lousto (2020, 2022); Jani et al. (2016); Hamilton et al. (2023) and is often computationally intensive, taking days to weeks to simulate a single BBH merger and its associated waveforms. In recent decades, significant advancements have been made in making NR codes more efficient in performing BBH simulations for systems where the masses of the two-component black holes are comparable, i.e., where is the mass ratio, is the mass of the primary black hole, and is the mass of the secondary black hole. While NR simulations have begun to push beyond the barrier Yoo et al. (2022); Lousto and Healy (2023); Boyle et al. (2019), long-duration simulations with high accuracy remain challenging and parameter space explorations mostly out of reach.
Alternatively, when , one can simplify the problem by assuming the secondary as a point-particle and solving the Teukolsky or the Regge-Wheeler-Zerilli (RWZ) equation Lopez-Aleman et al. (2003); Khanna (2004); Burko and Khanna (2007); Sundararajan et al. (2008, 2010); Zenginoglu and Khanna (2011); Fujita and Tagoshi (2004, 2005); Mano et al. (1996); Throwe (2010); O’Sullivan and Hughes (2014); Drasco and Hughes (2006) to obtain the far-field waveform. This theoretical framework is termed point-particle black-hole perturbation theory (ppBHPT) and is anticipated to yield accurate waveforms for binary systems characterized by sufficiently large mass ratios. Conversely, it is expected that the accuracy of ppBHPT waveforms will diminish as we move from the extreme mass ratio limit to the comparable mass regime.
Exploring the differences between waveforms derived from NR and ppBHPT in the comparable- and intermediate-mass-ratio regimes has long been a focal point of interest within the GW community Price and Pullin (1994); Lousto et al. (2010a, b); Nakano et al. (2011). This exploration serves a dual purpose. Firstly, it provides insights into the interaction between the NR and ppBHPT frameworks within the comparable mass regime, as well as the interplay between fully nonlinear and linear problem descriptions. Secondly, it aids in establishing the validity boundaries of the ppBHPT framework, paving the way for leveraging it to construct waveform models for BBH source characterization.
Two recent developments in extending the ppBHPT framework into the comparable mass regime are the development of a second-order self-force waveform model Pound and Wardell (2021); Miller and Pound (2021); Wardell et al. (2021) and the empirically observed efficacy of a simple model calibration technique, which we will refer to as - scaling Islam et al. (2022); Rifat et al. (2020). In this paper, we assess the applicability of the - scaling at the extreme end of the mass ratio regime: the equal mass binary. While previous studies explored the late-stage of equal mass collisions in great detail within the perturbative frameworks using the close-limit approximation Price and Pullin (1994); Anninos et al. (1995); Baker et al. (1997); Nollert et al. (1996); Khanna et al. (1999, 2000), we focus on the comparison between NR and ppBHPT for the full binary evolution.
The rest of this paper is organized as follows. In Section 2, we summarize our waveform data. Section 3 outlines various methods for determining the - parameters and demonstrates their effectiveness for an equal-mass system. Limitations and implications of our study are addressed in Section 4.
2 NR and ppBHPT data at
We generate our ppBHPT waveform using the BHPTNRSur1dq1e4 Islam et al. (2022) model 111Direct simulation using a time-domain Teukolsky solver fails in the equal-mass limit due to limitations in trajectory generation for the secondary black hole that incorporates adiabatic radiative corrections.. BHPTNRSur1dq1e4 is a reduced-order surrogate model trained on ppBHPT waveform data generated from a time-domain Teukolsky equation sourced by a test-particle whose adiabatic inspiral is driven by energy fluxes Sundararajan et al. (2007, 2008, 2010); Zenginoglu and Khanna (2011). This model is interfaced through the BHPTNRSurrogate’s package Field et al. available in the Black Hole Perturbation Toolkit BHP . While the model has been trained for mass ratios , surrogate models have previously demonstrated good performance when extrapolated beyond their training range Varma et al. (2019b); Islam et al. (2021). In our case, extrapolation to will bring in error, but these errors are expected to be small Islam et al. (2021).
3 Comparing NR and ppBHPT at
3.1 Model calibration setup
We now investigate the differences and connections between NR and ppBHPT at through the lens of the - scaling. The scaling reads Islam et al. (2022):
where and represents the NR and ppBHPT waveforms, respectively, as functions of the NR time, , and ppBHPT time, . The model calibration parameter values are found by solving an optimization problem
where the integral is taken over the last of NR data. This allows us to determine optimal values for and . To simplify the discussion, we focus only on the waveform’s mode and will use for brevity.
3.2 Setting a common mass scale
As discussed in Ref. Rifat et al. (2020); Islam et al. (2022), our Teukolsky solver sets the background black hole’s mass to and uses as ppBHPT’s mass scale, while the corresponding NR simulation sets the total mass to and uses as the mass scale. So before comparing waveforms, we should adjust the ppBHPT’s mass scale to use the NR convention. In the ppBHPT’s setup this would be , where upon setting , the mode should be adjusted according to the formula before comparing to NR. In Fig. 1, we show both the NR and ppBHPT waveforms after making this adjustment. While there are visual discrepancies, we find it remarkable that these waveforms are even somewhat similar, given that is well outside of the formal range of validity for point-particle perturbation theory.
While our - calibration procedure (Eq. (1)) automatically accounts for the different mass scales used, it also attempts to account for various physical effects missing in the ppBHPT setup. In keeping with conventions of previous work Rifat et al. (2020); Islam et al. (2022), we will apply the scaling procedure described in Sec. 3.1 to ppBHPT waveforms expressed units of . Because the mass scale is a freely specifiable choice for expressing dimensioned quantities in terms of dimensionless ones, changing the mass scale carries no more physical meaning than, for example, expressing time in seconds instead of hours. However, we suspect the numerical values of and do carry some physical meaning to account for missing physics (e.g. second-order self-force effects), although the interpretability of our phenomenological - scaling remains an open question. Yet, if the numerical values are to be interpreted as something physically relevant, the mass scale needs to be correctly accounted for. That is, one could imagine applying the scaling procedure described in Sec. 3.1 to ppBHPT waveforms expressed in terms of a mass scale resulting in a different set of parameters , which are related by and . For our system, and would be the relevant values to compute and where, for example, if , then NR and ppBHPT waveforms would be physically identical.
3.3 Effectiveness of the - scaling at
We first consider whether a simple scaling (Eq. 1) exists for equal-mass binaries. We find Eq. (2) is minimized at and where . In Fig. 1, we show both the NR and rescaled ppBHPT waveforms. Additionally, for comparison, we show the ppBHPT waveform after only accounting for the mass scale differences as described in Sec. 3.2.
It is worth noting that while the calibrated ppBHPT and NR waveforms exhibit excellent agreement during the inspiral phase, noticeable differences become apparent in the plunge-ringdown stage. These differences can be attributed to a combination of various simplifying assumptions in the ppBHPT framework; a non-inclusive list includes: (i) the final remnant black hole’s mass and spin value is assumed to be the same as the primary black hole’s initial mass and spin, (ii) the orbital plunge model is expected to be inaccurate in the comparable mass limit, (iii) finite size effects are likely to be critically important for an equal mass binary close to merger, and (iv) non-linear, strong gravity effects (e.g. second-order metric perturbations) may become important. For accurate modeling, these effects must be accounted for Islam and Khanna (2023); Islam et al. (2023). However, as extremely accurate GW models in the comparable mass regime already exist, even a sufficiently well-calibrated ppBHPT-based waveform model is unlikely to deliver competitive models at .
3.4 Exploring the time-dependence of - parameters from different methods
Clearly, the and parameters must have some temporal dependence. To illustrate this issue further, we compute time-varying and values using two different methods described in Ref. Islam and Khanna (2023). The first method seeks to find pointwise values of and such that the amplitude and phasing of the ppBHPT waveform match NR at each NR wave peak. We denote these estimates as and . For the second method, we selected the NR and ppBHPT data encompassing the final 35 cycles before the merger and segment them into seven consecutive 5-cycle windows. Each 5-cycle window is composed of 10 peaks. It is important to note that these waveform windows will have different time durations between NR and ppBHPT data but will maintain an identical number of peaks. Within each of these time windows, we minimize from Eq. (2) yielding local estimates of and applicable in each time window. We refer to these estimates as and . Using the second method, Figure 2 shows both the rescaled ppBHPT waveform (in red solid lines) and NR data (in black dashed lines) over the first and last 5-cycle time windows. Notice that in the last 5-cycle window, covering to , visible differences emerge between the rescaled ppBHPT waveform and NR, indicating the assumption of a time-independent - is no longer valid even within this restricted signal duration.
In Figure 3, we compare calibration parameter values obtained using all three methods described above. For comparison, we also show the scaling value of required to make the mass scale of ppBHPT and NR waveforms consistent, as discussed in Sec. 3.2. We find that and values obtained from these different approaches differ from , demonstrating that the calibration parameters are necessary to account for physical effects missing in the ppBHPT model. We also notice that and remain relatively constant in the early inspiral and then quickly rise near merger. This explains why the time-independent - scaling used in previous models Islam et al. (2022); Rifat et al. (2020) effectively matches the inspiral waveform but does not perform as well in the late inspiral-merger-ringdown phase. To emphasize the temporal evolution of the scaling parameters, we show derivatives of both and in Fig. 4.
3.5 Understanding - values at the ringdown
The final piece of our investigation focuses on the ringdown phase of the waveform where, as shown in Fig. 1, the most significant disparities between NR and rescaled ppBHPT waveforms are observed. This prompts the question of whether alternative and values can be found. We minimize Eq. (2) over the final of the signal finding optimal values and , which is significantly different from the inspiral-based values of and . In Figure 5, we show both the NR and rescaled ppBHPT ringdown signals using scaling parameter values that arise from different methods. We see that while the and parameters determined from the inspiral are not appropriate for the ringdown, a different set of parameter values – constant over the final M of the signal – can again deliver an accurate calibration.
4 Discussion & Conclusion
In this paper, we compared NR and ppBHPT waveforms for an equal mass binary black hole system. Given that is well outside of the formal range of validity for point-particle perturbation theory, this serves as a stress test for ppBHPT. We investigate the relationship between these ostensibly disparate frameworks through a phenomenological - scaling approach previously introduced as a simple yet effective method to calibrate ppBHPT waveforms to NR. This approach applied to the waveform’s (2,2) mode only requires two time-independent parameters to scale the overall amplitude and time coordinate.
Our findings indicate that the - scaling method Islam et al. (2022); Rifat et al. (2020) performs reasonably well for equal mass binaries, except during the merger-ringdown phase. This breakdown is likely associated with multiple limiting assumptions of ppBHPT, including the importance of finite size effects and the remnant blackhole’s mass and spin values. Furthermore, by generalizing our calibration parameters to be time-dependent, we illustrate that these parameters (i) have minimal temporal variation in the early inspiral and (ii) exhibit greater temporal variation as the binary system is about orbits before the merger. We also show that ringdown-specific scaling values – that are time-independent over the last of the signal – can be found. A comprehensive exploration of the merger-ringdown data for various mass ratios and spins will be important for future waveform calibration endeavors using this method. Despite these challenges, our investigation yields hopefully insightful observations regarding NR and ppBHPT waveforms under the constraints of equal mass binaries.
One potential limitation of our analysis is that the ppBHPT waveform is generated using BHPTNRSur1dq1e4 (a surrogate model) outside its training range. Yet direct simulation of a system with the ppBHPT framework was not possible as our time-domain Teukolsky solver Sundararajan et al. (2007, 2008, 2010); Zenginoglu and Khanna (2011) breaks down at the equal mass limit222More specifically, the trajectory generation for the secondary black hole that incorporates adiabatic radiative corrections fails in the equal-mass limit.. While surrogate models can be extrapolated a bit beyond their training range Varma et al. (2019b); Islam et al. (2021), this will bring in additional errors that might compromise studies that require very high accuracy.
Acknowledgements.
Submitted to a themed issue in honor of Prof. Jorge Pullin on his 60th anniversary. The authors acknowledge support of NSF Grants PHY-2106755, PHY-2307236 (G.K) and DMS-1912716, DMS-2309609 (T.I and G.K). SEF acknowledges support from NSF Grants PHY-2110496, DMS-2309609, and by UMass Dartmouth’s Marine and Undersea Technology (MUST) Research Program funded by the Office of Naval Research (ONR) under Grant No. N00014-23-1–2141. Simulations were performed on CARNiE at the Center for Scientific Computing and Data Science Research (CSCDR) of UMassD, which is supported by the ONR/DURIP Grant No. N00014181255 and the UMass-URI UNITY supercomputer supported by the Massachusetts Green High Performance Computing Center (MGHPCC).References
- Blackman et al. (2015) Blackman, J.; Field, S.E.; Galley, C.R.; Szilágyi, B.; Scheel, M.A.; Tiglio, M.; Hemberger, D.A. Fast and Accurate Prediction of Numerical Relativity Waveforms from Binary Black Hole Coalescences Using Surrogate Models. Phys. Rev. Lett. 2015, 115, 121102, [arXiv:gr-qc/1502.07758]. https://doi.org/10.1103/PhysRevLett.115.121102.
- Blackman et al. (2017a) Blackman, J.; Field, S.E.; Scheel, M.A.; Galley, C.R.; Ott, C.D.; Boyle, M.; Kidder, L.E.; Pfeiffer, H.P.; Szilágyi, B. Numerical relativity waveform surrogate model for generically precessing binary black hole mergers. Phys. Rev. 2017, D96, 024058, [arXiv:gr-qc/1705.07089]. https://doi.org/10.1103/PhysRevD.96.024058.
- Blackman et al. (2017b) Blackman, J.; Field, S.E.; Scheel, M.A.; Galley, C.R.; Hemberger, D.A.; Schmidt, P.; Smith, R. A Surrogate Model of Gravitational Waveforms from Numerical Relativity Simulations of Precessing Binary Black Hole Mergers. Phys. Rev. D 2017, 95, 104023, [arXiv:gr-qc/1701.00550]. https://doi.org/10.1103/PhysRevD.95.104023.
- Varma et al. (2019a) Varma, V.; Field, S.E.; Scheel, M.A.; Blackman, J.; Kidder, L.E.; Pfeiffer, H.P. Surrogate model of hybridized numerical relativity binary black hole waveforms. Phys. Rev. 2019, D99, 064045, [arXiv:gr-qc/1812.07865]. https://doi.org/10.1103/PhysRevD.99.064045.
- Varma et al. (2019b) Varma, V.; Field, S.E.; Scheel, M.A.; Blackman, J.; Gerosa, D.; Stein, L.C.; Kidder, L.E.; Pfeiffer, H.P. Surrogate models for precessing binary black hole simulations with unequal masses. Phys. Rev. Research. 2019, 1, 033015, [arXiv:gr-qc/1905.09300]. https://doi.org/10.1103/PhysRevResearch.1.033015.
- Islam et al. (2021) Islam, T.; Varma, V.; Lodman, J.; Field, S.E.; Khanna, G.; Scheel, M.A.; Pfeiffer, H.P.; Gerosa, D.; Kidder, L.E. Eccentric binary black hole surrogate models for the gravitational waveform and remnant properties: comparable mass, nonspinning case. Phys. Rev. D 2021, 103, 064022, [arXiv:gr-qc/2101.11798]. https://doi.org/10.1103/PhysRevD.103.064022.
- Bohé et al. (2017) Bohé, A.; et al. Improved effective-one-body model of spinning, nonprecessing binary black holes for the era of gravitational-wave astrophysics with advanced detectors. Phys. Rev. D 2017, 95, 044028, [arXiv:gr-qc/1611.03703]. https://doi.org/10.1103/PhysRevD.95.044028.
- Cotesta et al. (2018) Cotesta, R.; Buonanno, A.; Bohé, A.; Taracchini, A.; Hinder, I.; Ossokine, S. Enriching the Symphony of Gravitational Waves from Binary Black Holes by Tuning Higher Harmonics. Phys. Rev. D 2018, 98, 084028, [arXiv:gr-qc/1803.10701]. https://doi.org/10.1103/PhysRevD.98.084028.
- Cotesta et al. (2020) Cotesta, R.; Marsat, S.; Pürrer, M. Frequency domain reduced order model of aligned-spin effective-one-body waveforms with higher-order modes. Phys. Rev. D 2020, 101, 124040, [arXiv:gr-qc/2003.12079]. https://doi.org/10.1103/PhysRevD.101.124040.
- Pan et al. (2014) Pan, Y.; Buonanno, A.; Taracchini, A.; Kidder, L.E.; Mroué, A.H.; Pfeiffer, H.P.; Scheel, M.A.; Szilágyi, B. Inspiral-merger-ringdown waveforms of spinning, precessing black-hole binaries in the effective-one-body formalism. Phys. Rev. D 2014, 89, 084006, [arXiv:gr-qc/1307.6232]. https://doi.org/10.1103/PhysRevD.89.084006.
- Babak et al. (2017) Babak, S.; Taracchini, A.; Buonanno, A. Validating the effective-one-body model of spinning, precessing binary black holes against numerical relativity. Phys. Rev. D 2017, 95, 024010, [arXiv:gr-qc/1607.05661]. https://doi.org/10.1103/PhysRevD.95.024010.
- Husa et al. (2016) Husa, S.; Khan, S.; Hannam, M.; Pürrer, M.; Ohme, F.; Jiménez Forteza, X.; Bohé, A. Frequency-domain gravitational waves from nonprecessing black-hole binaries. I. New numerical waveforms and anatomy of the signal. Phys. Rev. D 2016, 93, 044006, [arXiv:gr-qc/1508.07250]. https://doi.org/10.1103/PhysRevD.93.044006.
- Khan et al. (2016) Khan, S.; Husa, S.; Hannam, M.; Ohme, F.; Pürrer, M.; Jiménez Forteza, X.; Bohé, A. Frequency-domain gravitational waves from nonprecessing black-hole binaries. II. A phenomenological model for the advanced detector era. Phys. Rev. D 2016, 93, 044007, [arXiv:gr-qc/1508.07253]. https://doi.org/10.1103/PhysRevD.93.044007.
- London et al. (2018) London, L.; Khan, S.; Fauchon-Jones, E.; García, C.; Hannam, M.; Husa, S.; Jiménez-Forteza, X.; Kalaghatgi, C.; Ohme, F.; Pannarale, F. First higher-multipole model of gravitational waves from spinning and coalescing black-hole binaries. Phys. Rev. Lett. 2018, 120, 161102, [arXiv:gr-qc/1708.00404]. https://doi.org/10.1103/PhysRevLett.120.161102.
- Khan et al. (2019) Khan, S.; Chatziioannou, K.; Hannam, M.; Ohme, F. Phenomenological model for the gravitational-wave signal from precessing binary black holes with two-spin effects. Phys. Rev. D 2019, 100, 024059, [arXiv:gr-qc/1809.10113]. https://doi.org/10.1103/PhysRevD.100.024059.
- Mroue et al. (2013) Mroue, A.H.; et al. Catalog of 174 Binary Black Hole Simulations for Gravitational Wave Astronomy. Phys. Rev. Lett. 2013, 111, 241104, [arXiv:gr-qc/1304.6077]. https://doi.org/10.1103/PhysRevLett.111.241104.
- Boyle et al. (2019) Boyle, M.; et al. The SXS Collaboration catalog of binary black hole simulations. Class. Quant. Grav. 2019, 36, 195006, [arXiv:gr-qc/1904.04831]. https://doi.org/10.1088/1361-6382/ab34e2.
- Healy et al. (2017) Healy, J.; Lousto, C.O.; Zlochower, Y.; Campanelli, M. The RIT binary black hole simulations catalog. Class. Quant. Grav. 2017, 34, 224001, [arXiv:gr-qc/1703.03423]. https://doi.org/10.1088/1361-6382/aa91b1.
- Healy et al. (2019) Healy, J.; Lousto, C.O.; Lange, J.; O’Shaughnessy, R.; Zlochower, Y.; Campanelli, M. Second RIT binary black hole simulations catalog and its application to gravitational waves parameter estimation. Phys. Rev. D 2019, 100, 024021, [arXiv:gr-qc/1901.02553]. https://doi.org/10.1103/PhysRevD.100.024021.
- Healy and Lousto (2020) Healy, J.; Lousto, C.O. Third RIT binary black hole simulations catalog. Phys. Rev. D 2020, 102, 104018, [arXiv:gr-qc/2007.07910]. https://doi.org/10.1103/PhysRevD.102.104018.
- Healy and Lousto (2022) Healy, J.; Lousto, C.O. Fourth RIT binary black hole simulations catalog: Extension to eccentric orbits. Phys. Rev. D 2022, 105, 124010, [arXiv:gr-qc/2202.00018]. https://doi.org/10.1103/PhysRevD.105.124010.
- Jani et al. (2016) Jani, K.; Healy, J.; Clark, J.A.; London, L.; Laguna, P.; Shoemaker, D. Georgia Tech Catalog of Gravitational Waveforms. Class. Quant. Grav. 2016, 33, 204001, [arXiv:gr-qc/1605.03204]. https://doi.org/10.1088/0264-9381/33/20/204001.
- Hamilton et al. (2023) Hamilton, E.; et al. A catalogue of precessing black-hole-binary numerical-relativity simulations 2023. [arXiv:gr-qc/2303.05419].
- Yoo et al. (2022) Yoo, J.; Varma, V.; Giesler, M.; Scheel, M.A.; Haster, C.J.; Pfeiffer, H.P.; Kidder, L.E.; Boyle, M. Targeted large mass ratio numerical relativity surrogate waveform model for GW190814. Physical Review D 2022, 106, 044001.
- Lousto and Healy (2023) Lousto, C.O.; Healy, J. Study of the intermediate mass ratio black hole binary merger up to 1000: 1 with numerical relativity. Classical and Quantum Gravity 2023, 40, 09LT01.
- Lopez-Aleman et al. (2003) Lopez-Aleman, R.; Khanna, G.; Pullin, J. Perturbative evolution of particle orbits around Kerr black holes: time domain calculation. Class. Quant. Grav. 2003, 20, 3259, [gr-qc/0303054]. https://doi.org/10.1088/0264-9381/20/14/320.
- Khanna (2004) Khanna, G. Teukolsky evolution of particle orbits around Kerr black holes in the time domain: elliptic and inclined orbits. Phys. Rev. D 2004, 69, 024016, [gr-qc/0309107]. https://doi.org/10.1103/PhysRevD.69.024016.
- Burko and Khanna (2007) Burko, L.; Khanna, G. Accurate time-domain gravitational waveforms for extreme-mass-ratio binaries. Europhys. Lett. 2007, 78, 60005, [gr-qc/0609002]. https://doi.org/10.1209/0295-5075/78/60005.
- Sundararajan et al. (2008) Sundararajan, P.A.; Khanna, G.; Hughes, S.A.; Drasco, S. Towards adiabatic waveforms for inspiral into Kerr black holes: II. Dynamical sources and generic orbits. Phys. Rev. D 2008, 78, 024022, [arXiv:gr-qc/0803.0317]. https://doi.org/10.1103/PhysRevD.78.024022.
- Sundararajan et al. (2010) Sundararajan, P.A.; Khanna, G.; Hughes, S.A. Binary black hole merger gravitational waves and recoil in the large mass ratio limit. Phys. Rev. D 2010, 81, 104009, [arXiv:gr-qc/1003.0485]. https://doi.org/10.1103/PhysRevD.81.104009.
- Zenginoglu and Khanna (2011) Zenginoglu, A.; Khanna, G. Null infinity waveforms from extreme-mass-ratio inspirals in Kerr spacetime. Phys. Rev. X 2011, 1, 021017, [arXiv:gr-qc/1108.1816]. https://doi.org/10.1103/PhysRevX.1.021017.
- Fujita and Tagoshi (2004) Fujita, R.; Tagoshi, H. New numerical methods to evaluate homogeneous solutions of the Teukolsky equation. Prog. Theor. Phys. 2004, 112, 415–450, [gr-qc/0410018]. https://doi.org/10.1143/PTP.112.415.
- Fujita and Tagoshi (2005) Fujita, R.; Tagoshi, H. New Numerical Methods to Evaluate Homogeneous Solutions of the Teukolsky Equation II. Solutions of the Continued Fraction Equation. Prog. Theor. Phys. 2005, 113, 1165–1182, [arXiv:gr-qc/0904.3818]. https://doi.org/10.1143/PTP.113.1165.
- Mano et al. (1996) Mano, S.; Suzuki, H.; Takasugi, E. Analytic solutions of the Teukolsky equation and their low frequency expansions. Prog. Theor. Phys. 1996, 95, 1079–1096, [gr-qc/9603020]. https://doi.org/10.1143/PTP.95.1079.
- Throwe (2010) Throwe, W.W.T. High precision calculation of generic extreme mass ratio inspirals. PhD thesis, Massachusetts Institute of Technology, 2010.
- O’Sullivan and Hughes (2014) O’Sullivan, S.; Hughes, S.A. Strong-field tidal distortions of rotating black holes: Formalism and results for circular, equatorial orbits. Phys. Rev. D 2014, 90, 124039, [arXiv:gr-qc/1407.6983]. [Erratum: Phys.Rev.D 91, 109901 (2015)], https://doi.org/10.1103/PhysRevD.91.109901.
- Drasco and Hughes (2006) Drasco, S.; Hughes, S.A. Gravitational wave snapshots of generic extreme mass ratio inspirals. Phys. Rev. D 2006, 73, 024027, [gr-qc/0509101]. [Erratum: Phys.Rev.D 88, 109905 (2013), Erratum: Phys.Rev.D 90, 109905 (2014)], https://doi.org/10.1103/PhysRevD.73.024027.
- Price and Pullin (1994) Price, R.H.; Pullin, J. Colliding black holes: The Close limit. Phys. Rev. Lett. 1994, 72, 3297–3300, [gr-qc/9402039]. https://doi.org/10.1103/PhysRevLett.72.3297.
- Lousto et al. (2010a) Lousto, C.O.; Nakano, H.; Zlochower, Y.; Campanelli, M. Intermediate Mass Ratio Black Hole Binaries: Numerical Relativity meets Perturbation Theory. Phys. Rev. Lett. 2010, 104, 211101, [arXiv:gr-qc/1001.2316]. https://doi.org/10.1103/PhysRevLett.104.211101.
- Lousto et al. (2010b) Lousto, C.O.; Nakano, H.; Zlochower, Y.; Campanelli, M. Intermediate-mass-ratio black hole binaries: Intertwining numerical and perturbative techniques. Phys. Rev. D 2010, 82, 104057, [arXiv:gr-qc/1008.4360]. https://doi.org/10.1103/PhysRevD.82.104057.
- Nakano et al. (2011) Nakano, H.; Zlochower, Y.; Lousto, C.O.; Campanelli, M. Intermediate-mass-ratio black hole binaries II: Modeling Trajectories and Gravitational Waveforms. Phys. Rev. D 2011, 84, 124006, [arXiv:gr-qc/1108.4421]. https://doi.org/10.1103/PhysRevD.84.124006.
- Pound and Wardell (2021) Pound, A.; Wardell, B. Black hole perturbation theory and gravitational self-force 2021. [arXiv:gr-qc/2101.04592].
- Miller and Pound (2021) Miller, J.; Pound, A. Two-timescale evolution of extreme-mass-ratio inspirals: waveform generation scheme for quasicircular orbits in Schwarzschild spacetime. Phys. Rev. D 2021, 103, 064048, [arXiv:gr-qc/2006.11263]. https://doi.org/10.1103/PhysRevD.103.064048.
- Wardell et al. (2021) Wardell, B.; Pound, A.; Warburton, N.; Miller, J.; Durkan, L.; Le Tiec, A. Gravitational waveforms for compact binaries from second-order self-force theory 2021. [arXiv:gr-qc/2112.12265].
- Islam et al. (2022) Islam, T.; Field, S.E.; Hughes, S.A.; Khanna, G.; Varma, V.; Giesler, M.; Scheel, M.A.; Kidder, L.E.; Pfeiffer, H.P. Surrogate model for gravitational wave signals from nonspinning, comparable-to large-mass-ratio black hole binaries built on black hole perturbation theory waveforms calibrated to numerical relativity. Phys. Rev. D 2022, 106, 104025, [arXiv:gr-qc/2204.01972]. https://doi.org/10.1103/PhysRevD.106.104025.
- Rifat et al. (2020) Rifat, N.E.M.; Field, S.E.; Khanna, G.; Varma, V. Surrogate model for gravitational wave signals from comparable and large-mass-ratio black hole binaries. Phys. Rev. D 2020, 101, 081502, [arXiv:gr-qc/1910.10473]. https://doi.org/10.1103/PhysRevD.101.081502.
- Anninos et al. (1995) Anninos, P.; Price, R.H.; Pullin, J.; Seidel, E.; Suen, W.M. Headon collision of two black holes: Comparison of different approaches. Phys. Rev. D 1995, 52, 4462–4480, [gr-qc/9505042]. https://doi.org/10.1103/PhysRevD.52.4462.
- Baker et al. (1997) Baker, J.G.; Abrahams, A.; Anninos, P.; Brandt, S.; Price, R.; Pullin, J.; Seidel, E. The Collision of boosted black holes. Phys. Rev. D 1997, 55, 829–834, [gr-qc/9608064]. https://doi.org/10.1103/PhysRevD.55.829.
- Nollert et al. (1996) Nollert, H.P.; Baker, J.G.; Price, R.; Pullin, J. A Slightly less grand challenge: Colliding black holes using perturbation techniques. In Proceedings of the 18th Texas Symposium on Relativistic Astrophysics, 10 1996, pp. 604–606, [gr-qc/9710011].
- Khanna et al. (1999) Khanna, G.; Baker, J.G.; Gleiser, R.J.; Laguna, P.; Nicasio, C.O.; Nollert, H.P.; Price, R.; Pullin, J. Inspiralling black holes: The Close limit. Phys. Rev. Lett. 1999, 83, 3581–3584, [gr-qc/9905081]. https://doi.org/10.1103/PhysRevLett.83.3581.
- Khanna et al. (2000) Khanna, G.; Gleiser, R.; Price, R.; Pullin, J. Close limit of grazing black hole collisions: Nonspinning holes. New J. Phys. 2000, 2, 3, [gr-qc/0003003]. https://doi.org/10.1088/1367-2630/2/1/303.
- Sundararajan et al. (2007) Sundararajan, P.A.; Khanna, G.; Hughes, S.A. Towards adiabatic waveforms for inspiral into Kerr black holes. I. A New model of the source for the time domain perturbation equation. Phys. Rev. D 2007, 76, 104005, [gr-qc/0703028]. https://doi.org/10.1103/PhysRevD.76.104005.
- (53) Field, S.; Islam, T.; Khanna, G.; Rifat, N.; Varma, V. BHPTNRSurrogate. http://bhptoolkit.org/BHPTNRSurrogate/.
- (54) Black Hole Perturbation Toolkit. (bhptoolkit.org).
- Islam and Khanna (2023) Islam, T.; Khanna, G. Interplay between numerical relativity and perturbation theory: Finite size effects. Phys. Rev. D 2023, 108, 044012, [arXiv:gr-qc/2306.08767]. https://doi.org/10.1103/PhysRevD.108.044012.
- Islam et al. (2023) Islam, T.; Field, S.E.; Khanna, G. Remnant black hole properties from numerical-relativity-informed perturbation theory and implications for waveform modelling 2023. [arXiv:gr-qc/2301.07215].
- Islam and Khanna (2023) Islam, T.; Khanna, G. On the approximate relation between black-hole perturbation theory and numerical relativity 2023. [arXiv:gr-qc/2307.03155].